# Answer When Needed, Forget When Not: Language Models Pretend to Forget via In-Context Knowledge Unlearning

### v1

@@ Abstract
As large language models (LLMs) are applied across diverse domains, the ability to selectively unlearn specific information has become increasingly essential. For instance, LLMs are expected to provide confidential information to authorized internal users, such as employees or trusted partners, while withholding it from external users, including the general public and unauthorized entities. In response to this challenge, we propose a novel method termed “in-context knowledge unlearning”, which enables the model to selectively forget information in test-time based on the context of the query. Our method fine-tunes pre-trained LLMs to enable prompt unlearning of target knowledge within the context, while preserving other knowledge. Experiments on the TOFU and AGE datasets using Llama2-7B/13B and Mistral-7B models show our method achieves up to 95% forgetting accuracy while retaining 80% of unrelated knowledge, significantly outperforming baselines in both in-domain and out-of-domain scenarios. Further investigation into the model’s internal behavior revealed that while fine-tuned LLMs generate correct predictions in the middle layers and maintain them up to the final layer, they make the decision to forget at the last layer, i.e., “LLMs pretend to forget”. Our findings offer valuable insights into enhancing the robustness of unlearning mechanisms in LLMs, setting a foundation for future research in the field.  [^1] Code is available at https://anonymous.4open.science/r/test-time-in-context-unlearning
@@
Figure 1: Method overview. (1) Without unlearning, LLMs output any answers to given inputs. (2) Some prior unlearning methods (e.g.Pawelczyk et al. (2023)) unlearn a specific knowledge and cause hallucination. (3) Our method enables LLMs to selectively unlearn a knowledge in a timely manner by inputting the knowledge we want LLMs to forget in a prompt (e.g., <<UNL>>Paris<</UNL>>). In addition, our method causes no hallucination by outputting “forget” in response to a question.
@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4 OpenAI et al. (2024), have significantly transformed various sectors by providing advanced capabilities in information processing and generation.
@@
The widespread deployment of such models, however, introduces complex challenges related to privacy and the ethical use of information. Particularly, the indiscriminate supply of sensitive or domain-specific information by LLMs raises significant concerns, necessitating mechanisms for selective information handling based on the users’ context Das et al. (2024). To enhance privacy and ethical use of LLMs, prior works have explored several approaches, including differential privacy Abadi et al. (2016), federated learning Geyer et al. (2018), and knowledge distillation Jiang et al. (2023b). Despite their contributions, these methods often compromise between maintaining privacy and sustaining model performance.
@@
The concept of “test-time adaptation” Liang et al. (2023) or “in-context learning” Dong et al. (2024) offers a dynamic approach to model adaptation, yet it fails to adequately address the selective forgetting of sensitive information. For example, an LLM used within a corporate environment to streamline project management needs to retain substantial industry-specific knowledge while being capable of “forgetting” proprietary company data or sensitive information when accessed by unauthorized external consultants. This scenario underscores the critical need for a mechanism that enables LLMs to selectively forget or withhold sensitive information based on the query context without compromising their overall utility and performance.
@@
This paper introduces “in-context knowledge unlearning”, a novel approach designed to equip LLMs with the capability of selective forgetting in test-time, based on the query context. The overview of our method is given in Figure 1. We develop unlearning tokens that, when applied during inference, enable the model to selectively ignore information pertaining to specified domains. Through comprehensive experimentation, we validate the efficacy of our approach in facilitating domain-specific unlearning without compromising the model’s general performance. Specifically, we conducted experiments on the TOFU and AGE datasets using Llama2-7B/13B and Mistral-7B models, showing that our method achieves up to 95% forgetting accuracy while retaining 80% of unrelated knowledge, significantly outperforming baselines in both in-domain and out-of-domain scenarios.
@@
Moreover, further investigations into the model’s internal behavior revealed that while fine-tuned LLMs generate correct predictions in the middle layers and maintain them up to the final layer, they make the decision to forget at the last layer, i.e., “LLMs pretend to forget”. This finding not only enriches our understanding of selective information handling in LLMs but also sets a foundation for future research to improve the robustness of models across sensitive and regulated domains.
@@
Method Test-Time Unlearning Knowldge Unlearning Non-Hallucination Output
Gradient Ascent Golatkar et al. (2020) × × ×
@@ 2 Related Work
Our method leverages in-context learning (ICL) for knowledge unlearning. ICL enables LLMs to adapt to new tasks flexibly by incorporating data into the context of input sequence, rather than fine-tuning, which explicitly updates weights Brown et al. (2020); Dong et al. (2023); Liu et al. (2023). Exploring the full capabilities of ICL remains an active area of research, with recent studies empirically investigating its potential by examining in-context example design Garg et al. (2022); Liu et al. (2022); Min et al. (2022); Liu et al. (2023).
Pawelczyk et al. (2023) explored methods for performing in-context unlearning. This study focuses on text classification tasks where the labels of specific instances are flipped to facilitate in-context unlearning. However, this approach has limitations as it primarily assesses unlearning in terms of text classification ability rather than actual knowledge. Furthermore, the method trains the model to generate incorrect outputs, which is not meant true forgetting.
In contrast, our study introduces unique characteristics that address these issues. We specifically investigate knowledge unlearning within an in-context learning framework. Moreover, by defining unlearning as the ability to “forget” we ensure that our approach avoids merely generating errors or irrelevant information, thereby achieving a more effective and appropriate form of unlearning.
@@
Table 1 compares our method with existing unlearning techniques. Test-time unlearning refers to the process of selectively removing a specific concept or knowledge from a trained model. Knowledge unlearning refers to forgetting world knowledge, e.g., “The capital of France is Paris”.
For example, Gradient Ascent Golatkar et al. (2020) lacks test-time unlearning and only removes global knowledge. ROME Meng et al. (2022) and Knowledge Sanitization Ishibashi and Shimodaira (2024) require separate training to unlearn specific knowledges so that these methods cannot perform test-time unlearning. While ICUL (In-context Unlearning) Pawelczyk et al. (2023) achieves test-time unlearning, it merely changes a ground-truth label or word of target instance within the in-context prompt, so this approach inevitably outputs hallucinations.
Unlike existing methods, our approach achieves test-time unlearning, knowledge unlearning, and non-hallucination output simultaneously. In other words, our approach addresses the prior limitations and offers a comprehensive solution for selective forgetting.
@@ 3.1 In-context Knowledge Unlearning
In the context of in-context knowledge unlearning, a pre-trained auto-regressive language model modifies its response to a query $q$ by disregarding specific undesired information $u$. The response $r$ is generated according to the conditional probability distribution:
@@ 3.2 Unlearning Tokens
We introduce unlearning tokens to enable selective forgetting in LLMs during inference. These tokens are implemented by encapsulating the target information $u$ with <<UNL>> and <</UNL>>. For example, to forget ‘Paris’, the input would be: <<UNL>>Paris<</UNL>>. This corresponds to the information to be forgotten, $u$ in the Equation 1. The model is instructed to disregard the enclosed information during processing, effectively modifying its output distribution $P_{\theta}$. To integrate these tokens, we fine-tune the model using methods such as Low Rank Adaptation (LoRA), full model fine-tuning, or other parameter-efficient fine-tuning (PEFT) techniques, adjusting $\theta$ to recognize and respond to the unlearning tokens.
@@ 3.3 Loss Function
The loss function for our in-context knowledge unlearning method is designed to selectively suppress specific information while retaining other useful knowledge. This loss function consists of two main components: $L_{forget}$ and $L_{retain}$.
@@
where $r_{i}$ represents the tokens in the response to a given query.
@@
By minimizing this loss function, the model learns to balance the ability to selectively "forget" specified information with the ability to retain other useful knowledge. This approach enables the LLM to manage information appropriately based on the context, effectively implementing in-context knowledge unlearning.
@@ 4.1 Models
  • Llama2-7B/13B Touvron et al. (2023): Llama 2 is a family of large language models (LLMs) developed by Meta. Llama 2-7B and Llama 2-13B are two variants with 7 billion and 13 billion parameters, respectively. These models exhibit strong performance across a wide range of natural language processing tasks, making them suitable for tasks such as text generation, summarization, and translation. Llama2-7B and Llama2-13B specifically utilize the chat model configuration.
  • Mistral-7B Jiang et al. (2023a): Mistral-7B is an open-source LLM with 7 billion parameters developed by Mistral AI. This model is known for its high performance and low resource requirements, making it an attractive option for developers with limited resources. Mistral-7B has demonstrated performance comparable to other open-source LLMs on a variety of language processing tasks and employs the instruct model configuration.
@@ 4.2 Datasets
  • Age Dataset Annamoradnejad and Annamoradnejad (2022): The Original Age dataset contains structured information about the life, work, and death of over 1 million deceased famous individuals. From this, 180 individuals are randomly sampled, and a set of 5 questions and answers (QAs) is created for each individual. This dataset is employed to further investigate the models’ ability to generalize selective forgetting across various contexts. It includes 600 training samples and 300 validation samples.
@@ 4.3 Compared Methods
In this paper, we compare our proposed method with four other approaches capable of test-time unlearning:
@@
  • Zero-shot Prompting: This method as our baseline evaluation for in-context knowledge unlearning using a hard prompt. The model is directly instructed to disregard certain information specified within the prompt, providing a clear basis for comparison with more sophisticated unlearning methods. The specific prompt format used to guide the model’s behavior regarding memory retention and deletion is illustrated in Figure 2.
@@
  • Few-shot Prompting: This method builds upon the zero-shot approach by incorporating examples from the training data. In addition to the format shown in Figure 2, we randomly select and include five samples from the training data in the prompt.It is selected so that at least one data sample to be forgotten is included.
@@
  • Gradient Ascent: This method applies gradient ascent on the data to be forgotten and gradient descent on the data to be retained. To enable test-time unlearning, we incorporate the <<UNL>> token during training.
@@
  • ICUL (In-context Unlearning): This approach adds the data to be forgotten at test-time to the context, along with several instances of data to be retained. For the data to be forgotten, the answer is replaced with a randomly selected response from the training data.
@@
These methods provide a comprehensive comparison framework, allowing us to evaluate the effectiveness of our proposed in-context knowledge unlearning technique against established and emerging approaches in the field.
<s>[INST]
<<SYS>>
## Rules
- If there is no <<UNL>> token, please act normally.
<</SYS>>
<<UNL>>{unlearning word}<</UNL>>{query}
[/INST]
Figure 2: Zero Shot Prompt
4.4 Evaluation
To assess the effectiveness of our “in-context knowledge unlearning” method, we employ two primary metrics:
  • Forget: The proportion of instances where the model outputs “forgot”. A higher score indicates that the model is effectively “forgetting” the instructed information. Unlike previous studies Ishibashi and Shimodaira (2024), this metric directly assesses the model’s ability to acknowledge its intentional forgetting.
  • Retain: The proportion of questions the model answers correctly. A higher score suggests that the model is maintaining its essential knowledge.
These metrics were evaluated under two scenarios:
  • In-domain: The learning data was split into training and validation sets in an 8:2 ratio.
  • Out-of-domain: Evaluation was conducted using the world facts data from the TOFU dataset.
5 Result
5.1 Performance Results
Table 2 shows the results of our experiments across various unlearning methods and tasks. Our proposed method consistently outperforms baseline approaches for both LLaMA2 and Mistral models. For LLaMA2 (7B), we achieve ‘Forget’ and ‘Retain’ scores of 85.0% and 80.0%, respectively for in-domain data, significantly surpassing the zero-shot baseline. Out-of-domain performance remains strong with 92.3% ‘Forget’ and 42.7% ‘Retain’ scores. LLaMA2 (13B) shows even better results, particularly for in-domain scenarios, with perfect ‘Forget’ scores (100.0%) and high ‘Retain’ scores (80.0%).
Our method maintains competitive performance on standard NLP tasks such as BoolQ, HellaSwag, and WinoGrande, with minimal degradation compared to baseline models. This suggests that the unlearning process does not significantly impact the model’s general language understanding capabilities. Compared to other unlearning methods like Few-Shot Prompting, Gradient Ascent, and In-Context Unlearning, our approach consistently achieves a better balance between forgetting targeted information and retaining general knowledge.
These results demonstrate the effectiveness of our in-context knowledge unlearning method in enabling large language models to selectively forget information while maintaining overall performance across various NLP tasks.
@@ 5.1 Performance Results
Table 2: Comparison of various unlearning methods across multiple tasks. The table includes ‘Forget’ and ‘Retain’ scores for the TOFU dataset in both in-domain and out-of-domain scenarios, as well as performance metrics for additional tasks such as BoolQ, HellaSwag, WinoGrande, ARC-e, ARC-c, OBQA, and RACE-high.

### v2

@@ Abstract
As large language models (LLMs) are applied across diverse domains, the ability to selectively unlearn specific information is becoming increasingly essential. For instance, LLMs are expected to selectively provide confidential information to authorized internal users, such as employees or trusted partners, while withholding it from external users, including the general public and unauthorized entities. Therefore, we propose a novel method termed “in-context knowledge unlearning”, which enables the model to selectively forget information in test-time based on the query context. Our method fine-tunes pre-trained LLMs to enable prompt unlearning of target knowledge within the context, while preserving unrelated information. Experiments on TOFU, AGE and RWKU datasets using Llama2-7B/13B and Mistral-7B models demonstrate that our method achieves up to 95% forget accuracy while retaining 80% of unrelated knowledge, significantly outperforming baselines in both in-domain and out-of-domain scenarios. Further investigation of the model’s internal behavior revealed that while fine-tuned LLMs generate correct predictions in the middle layers and preserve them up to the final layer. However, the decision to forget is made only at the last layer, i.e. “LLMs pretend to forget”. Our findings offer valuable insight into the improvement of the robustness of the unlearning mechanisms in LLMs, laying a foundation for future research in the field.  [^1] Code is available at https://github.com/seele1917/test-time-in-context-unlearning
@@
Figure 1: Method overview. (1) Without unlearning, LLMs output any answers to specific inputs. (2) Certain prior unlearning methods (e.g.,Pawelczyk et al. (2023)) attempt to unlearn specific knowledge but may cause hallucinations. (3) Our method enables LLMs to selectively unlearn knowledge in a timely manner by inputting the knowledge we want LLMs to forget in a prompt (e.g., <<UNL>>Paris<</UNL>>). In contrast to In-context Unlearning (ICUL)  Pawelczyk et al. (2023), our method causes no hallucination by outputting “forget” in response to a question.
@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4 OpenAI et al. (2024), have significantly transformed various sectors by providing advanced capabilities in information processing and text generation.
@@
However, the widespread deployment of such models introduces complex challenges related to privacy and the ethical use of information. In particular, the indiscriminate provision of sensitive or domain-specific information by LLMs raises significant concerns, necessitating mechanisms for selective information handling based on the user context Das et al. (2024). To improve the privacy and ethical use of LLMs, previous works have explored several approaches, including differential privacy Abadi et al. (2016), federated learning Geyer et al. (2018), and knowledge distillation Jiang et al. (2023b). Despite their contributions, these methods often compromise between privacy and model performance.
@@
The concept of “test-time adaptation” Liang et al. (2023) or “in-context learning” Dong et al. (2024) offers a dynamic approach to model adaptation, yet it fails to adequately address selective forgetting of sensitive information. For example, an LLM deployed in a corporate environment to streamline project management should retain substantial industry-specific knowledge while selectively “forgetting” proprietary company data or sensitive information when accessed by unauthorized external consultants. This scenario underscores the critical need for a mechanism that enables LLMs to selectively forget or withhold sensitive information based on the query context without compromising their overall utility and performance.
@@
This paper introduces “in-context knowledge unlearning”, a novel approach that enables LLMs to selectively forget information at test-time based on the query context. The overview of our method is provided in Figure 1. We developed unlearning tokens that, when applied during inference, enable the model to selectively ignore information pertaining to specified domains. Through comprehensive experimentation, we validated the efficacy of our approach in facilitating domain-specific unlearning without compromising the general performance of the model. Specifically, we conducted experiments on the TOFU, AGE, and RWKU datasets Maini et al. (2024); Annamoradnejad and Annamoradnejad (2022); Jin et al. (2024) using Llama2-7B/13B and Mistral-7B models, showing that our method achieves up to 95% forget accuracy while retaining 80% unrelated knowledge, significantly outperforming baselines in both in-domain and out-of-domain scenarios.
@@
Moreover, further investigations into the model’s internal behavior revealed that while fine-tuned LLMs generate correct predictions in the middle layers and sustain them through to the final layer, the decision to forget occurs only at the last layer, i.e., “LLMs pretend to forget”. This finding enriches our understanding of selective information handling in LLMs and lays a foundation for future research to improve the robustness of models across sensitive and regulated domains.
@@
Method Test-Time Unlearning Non-Hallucination Output
Gradient Ascent Golatkar et al. (2020) × ×
@@ 2 Related Work
Our method leverages in-context learning (ICL) for knowledge unlearning. ICL enables LLMs to adapt flexibly to new tasks by incorporating data within the input sequence context, rather than through weight updates as in fine-tuning Brown et al. (2020a); Dong et al. (2023); Liu et al. (2023). Exploring the full capabilities of ICL remains an active area of research, with recent studies empirically investigating its potential by examining in-context example design Garg et al. (2022); Liu et al. (2022); Min et al. (2022); Liu et al. (2023).
Pawelczyk et al. (2023) explored methods for performing in-context unlearning. This study focuses on text classification tasks where the labels of specific instances are flipped to facilitate in-context unlearning. However, this approach has limitations as it primarily assesses unlearning in terms of text classification ability rather than actual knowledge. Furthermore, the method trains the model to generate incorrect outputs, which does not constitute true forgetting.
In contrast, our study introduces unique characteristics that address these issues. We specifically investigated knowledge unlearning within an in-context learning framework. Moreover, by defining unlearning as the ability to “forget,” we ensure that our approach avoids merely generating errors or irrelevant information, thereby achieving an effective and appropriate form of unlearning.
@@
Table 1 compares our method with existing unlearning techniques. The column of “Test-Time Unlearning” refers to selectively removing specific concepts or knowledge from a trained model during inference depending on the input context. “Non-Hallucination Output” refers to outputing "forgot" instead of producing a hallucinated answer.
Gradient Ascent Golatkar et al. (2020) lacks test-time unlearning and erases only the specific global knowledge during training. Additionally, because it modifies model parameters in a disruptive manner, it can cause hallucinated outputs.
Certain methods such as ROME Meng et al. (2022), Knowledge Sanitization Ishibashi and Shimodaira (2024), and Preference-based Unlearning Maini et al. (2024); Zhang et al. (2024) require dedicated retraining to remove specific knowledge, making them incapable of test-time unlearning. However, these approaches explicitly train the model to respond with "forgot", ensuring that it does not produce hallucinations.
While ICUL (In-Context Unlearning) Pawelczyk et al. (2023) achieves test-time unlearning, it operates by simply altering labels or tokens within the prompt context, which inherently induces hallucinated outputs.
In contrast to these existing methods, our approach achieves test-time unlearning, knowledge unlearning, and non-hallucinated outputs simultaneously, addressing prior limitations and offering a comprehensive solution for selective forgetting.
@@ 3.1 In-context Knowledge Unlearning
In the context of in-context knowledge unlearning, a pretrained auto-regressive language model modifies its response to a query $q$ by disregarding specific undesired information $u$. The response $r$ is generated according to the conditional probability distribution:
@@ 3.2 Unlearning Tokens
We introduce unlearning tokens to enable selective forgetting in LLMs during inference. These tokens are implemented by encapsulating the target information $u$ with <<UNL>> and <</UNL>>. For example, to forget ‘Paris’, the input will be: <<UNL>>Paris<</UNL>>. This corresponds to the information to be forgotten $u$ in Equation 1. The model is instructed to ignore the enclosed information during processing, effectively modifying its output distribution $P_{\theta}$. To integrate these tokens, we fine-tune the model using methods, such as Low-Rank Adaptation (LoRA), full model fine-tuning, or other parameter-efficient fine-tuning (PEFT) techniques, adjusting $\theta$ to recognize and respond to the unlearning tokens.
@@ 3.3 Loss Function
The loss function for our in-context knowledge unlearning method is designed to selectively suppress specific information while preserving other useful knowledge. This loss function consists of two main components: $L_{forget}$ and $L_{retain}$.
@@
where $r_{i}$ represents the tokens in the response to a specific query.
@@
By minimizing this loss function, the model learns to balance selective "forget" of specified information with the retention of other valuable knowledge. This approach enables the LLMs to manage information appropriately based on the context, effectively implementing in-context knowledge unlearning.
@@ 4.1 Models
  • Llama2-7B/13B Touvron et al. (2023): Llama 2 is a family of large language models (LLMs) developed by Meta. Llama 2-7B and Llama 2-13B are two variants with 7 billion and 13 billion parameters, respectively. These models exhibit strong performance on a wide range of natural language processing tasks, making them suitable for tasks, such as text generation, summarization, and translation. We employ chat configurations for Llama2-7B and Llama2-13B.
  • Mistral-7B Jiang et al. (2023a): Mistral-7B is an open-source LLM with 7 billion parameters developed by Mistral AI. This model is recognized for its high performance and low resource requirements, making it an attractive option for developers with limited resources. Mistral-7B has demonstrated performance comparable to other open-source LLMs on a variety of language processing tasks and employs the instruct model configuration.
@@ 4.2 Datasets
  • Age Dataset Annamoradnejad and Annamoradnejad (2022): The Original Age dataset contains structured information about the life, work, and death of over 1 million deceased famous individuals. Accordingly, 180 individuals are randomly sampled, and a set of five question-answer (QAs) pairs generated for each individual. This dataset is employed to further investigate the models’ ability to generalize selective forgetting across various contexts. It includes 600 training samples and 300 validation samples.
  • RWKU Dataset Jin et al. (2024): The Real-World Knowledge Unlearning (RWKU) dataset is a benchmark specifically designed for large language models (LLMs) to assess their ability to unlearn specific knowledge. It contains 200 real-world unlearning targets and 13,131 multilevel forget probes, including 3,268 fill-in-the-blank probes, 2,879 question-answer probes, and 6,984 adversarial-attack probes. In our experiments, we employed 20% of the question-answer data as out-of-domain data to evaluate the models’ performance in unlearning specific knowledge while maintaining overall functionality.
@@ 4.3 Compared Methods
In this paper, we compare our proposed method with four other approaches that can test-time unlearning:
@@
  • Zero-shot Prompting: We used this method as our baseline for evaluating in-context knowledge unlearning using a hard prompt. The model is directly instructed to disregard certain information specified within the prompt, providing a clear basis for comparison with more sophisticated unlearning methods. The specific prompt format used to guide the model’s behavior regarding memory retention and deletion is illustrated in Figure 5 of Appendix E.
@@
  • Few-shot Prompting Brown et al. (2020b): This method builds on the zero-shot approach by incorporating examples from the training data. In addition to the format shown in Figure 5, we randomly selected and included five samples from the training data in the prompt such that at least one data sample to be forgotten was included. Detailed examples of the few-shot prompts used can be found in Figure 6 of the Appendix E.
@@
  • Gradient Ascent Golatkar et al. (2020): This method applies gradient ascent to the data to be forgotten and gradient descent to the data to be retained. To enable test-time unlearning, we incorporated the <<UNL>> token during training.
@@
  • ICUL (In-context Unlearning) Pawelczyk et al. (2023): This approach adds the data to be forgotten at test-time to the context, along with several instances of data to be retained. In case the data are forgotten, the answer is replaced with a randomly selected response from the training data. Detailed examples of the ICUL prompts used are presented in Figure 7 of Appendix E
@@
These methods provide a comprehensive comparison framework for evaluating the effectiveness of our proposed in-context knowledge unlearning technique against established and emerging approaches in the field.
@@
Table 2: Comparison of various unlearning methods across multiple tasks. The table includes ‘Forget’ and ‘Retain’ scores for the TOFU dataset in both in-domain and out-of-domain scenarios, as well as performance metrics for additional tasks, such as BoolQ, HellaSwag, WinoGrande, ARC-e, ARC-c, OBQA, and RACE-high.
@@
4.4 Evaluation
To assess the effectiveness of our “in-context knowledge unlearning” method, we employed two primary metrics:
  • Forget: The proportion of instances where the model correctly outputs “forgot” if the knowledge asked in the question matches the knowledge to be forgotten set in the in-context. A higher score indicates that the model is effectively “forgetting” the instructed information. Contrarily to previous studies Ishibashi and Shimodaira (2024), this metric directly assesses the ability of the model to acknowledge its intentional forgetting. If the model outputs anything else, the answer is treated as a hallucination.
  • Retain: The proportion of questions the model correctly answers if the knowledge asked in the question does not match the knowledge to be forgotten set in the in-context. A higher score suggests that the model is maintaining its essential knowledge.
Note that ToFU’s “Forgetting Quality” and “Model Utility” metrics are not used here because they assume free-form answers, whereas our model returns only forgot token or the factual answer. Nevertheless, the same aspects are captured by our Forget and Retain accuracies.
These metrics were evaluated in two scenarios:
  • In-domain: The learning data (TOFU Real Authors, Age dataset, and RWKU) was divided into training and test sets using an 8:2 ratio. We evaluated the performance of models on the test sets after training them on the train sets.
  • Out-of-domain: We evaluated on the world facts data from the TOFU dataset and RWKU dataset.
In both scenarios, neither the in-domain nor the out-of-domain test queries include any tokens marked for unlearning during training (i.e., the «UNL»-enclosed terms). Consequently, simply memorizing these tokens to respond "forgot" will worsen the evaluation results. However, the model must rely purely on input context at test-time to flexibly detect knowledge to be forgotten, which enables the evaluation of true context-dependent forgetting ability.
5 Result
5.1 Performance Results
Table 2 shows the results of our experiments in various unlearning methods and tasks. Our proposed method consistently outperforms baseline approaches for both LLaMA2 and Mistral models.
For LLaMA2 (7B), we achieve ‘Forget’ and ‘Retain’ scores of 85.0% and 80.0%, respectively for in-domain data, significantly surpassing the zero-shot baseline. Out-of-domain performance remains strong with 92.3% ‘Forget’ and 42.7% ‘Retain’ scores. Notably, our out-of-domain evaluations are conducted using TOFU’s world facts dataset, and additional results obtained with RWKU are provided in appendix D. LLaMA2 (13B) shows better results, particularly for in-domain scenarios, with perfect ‘Forget’ scores (100.0%) and high ‘Retain’ scores (80.0%).
[…]
These results demonstrate the effectiveness of our in-context knowledge-unlearning method, enabling large language models to selectively forget information while maintaining overall performance across various NLP tasks.